Hydrogen gas can be produced by the reaction of hydrochloric acid with aluminum metal. How many grams of aluminum would be needed to fill-up a 250-mL container with hydrogen gas if the density of hydrogen gas is 0.84 g/L if the yield is only 75%.


1
Expert's answer
2021-01-11T03:53:25-0500

3HCl + Al gives AlCl3 + 1.5H2

density = mass/volume

0.84g/L = mass/0.25L

mass = 0.84×0.25 = 0.2125g

Percentage yield = "\frac{obtained yield}{theoreticalyield}" × 100

75 = "\frac{0.2125}{theoreticalyield}" × 100

theoretical yield = (100×0.2125)/75 = 0.2833g

mass of aluminum required = 0.2833/1.5

= 0.1888g